OpenStudy (anonymous):

Determine which of the lines, if any, are parallel Line a: 4y + x = 8 Line b: 2y + x = 4 Line c: 2y = -3x + 6 please help

OpenStudy (anonymous):

a and b are parallel

OpenStudy (anonymous):

Thank you, but do you think you could explain why?

OpenStudy (anonymous):

yeah sure: they have the same slope; therefore, the lines will be parallel

OpenStudy (anonymous):

oh okay, thank you so much! :)

OpenStudy (anonymous):

no problem. so you understand why now?

OpenStudy (anonymous):

yeah but can I ask you one more question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The next one I have is Line a: 3y-x=6 Line b: 3y=x+18 Line c: 3y-2x=9 I think I just don't know how to get the slope

OpenStudy (anonymous):

y=mx+b have you seen this equation before ?

OpenStudy (anonymous):

yeah my math teacher showed me in class

OpenStudy (anonymous):

okay so the m next to x is known as the slope

OpenStudy (anonymous):

so get your equations into that form and check to see if the m matches for your equations

OpenStudy (anonymous):

if the the m for equation 1 is the same as the m for equation 2, then that means both eqn 1 and 2 have the same slope, and if the slopes are the same, the lines will be parallel

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

A little bit

OpenStudy (anonymous):

I did some of this is class yesterday but the problems that I was doing were different and now im doing my homework so I was doing problems that were like The line passes through (-1, 3) and (1,9) and for those problems I used the m thing but then I Got to these problems and now I'm confused...

OpenStudy (anonymous):

just remember that m = slope if two lines have the same m, then those lines will be parallel. if i could draw with my mouse I would show you, but I can't XD

OpenStudy (anonymous):

XD It's okay. Thank you :) But do you know what this is Line a: 3y-x=6 Line b: 3y=x+18 Line c: 3y-2x=9

OpenStudy (anonymous):

line a: 3y-x=6 ---> y=x/3 +6/3 ----> y=x/3 +3 line b: 3y=x+18 -->y=x/3 +18/3 ---> y=x/3 +6 line c: 3y-2x=9 --> y=2x/3+9/3 ---->y=2x/3 +3 which slope match?

OpenStudy (anonymous):

a and c?

OpenStudy (anonymous):

a and b right? because m= 1/3 for a and b. for c, m=2/3

OpenStudy (anonymous):

ohh now I get it

OpenStudy (anonymous):

okay good. :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

no problem :)